# Schemex: Interactive Structural Abstraction from Examples with Contrastive Refinement

### v1

@@ 3. Schemex System
Schemex (see Figure 2) is an interactive visual workflow that assists users in transforming a set of examples into actionable schemas through three AI-assisted stages: clustering, abstraction, and refinement via contrasting examples. If the input examples are multimodal, the system first performs a preprocessing step that converts them into structured textual representations using models such as GPT-4V and Whisper. Next, the workflow clusters the examples into distinct groups based on their latent similarities. For each cluster, the system examines the examples to derive key dimensions and dimension-specific attributes. Overall attributes of the examples within the cluster are also inferred, forming an initial schema comprising both dimensions and attributes. Finally, Schemex refines the schema through contrastive refinement: it generates outputs based on the initial schema, compares them to the original human-authored examples, and iteratively improves the schema. Throughout the workflow, each intermediate result is presented in an interactive node. Clicking a node reveals a side panel containing more detailed information, such as definitions and example mapping matrices, helping users effectively ground their schema inspection and understanding in the example set.
Schemex is implemented using the Flask/React Flow web framework. It is powered by a reasoning LLM (o3-mini-high) for reasoning and analysis tasks, and a general-purpose LLM (GPT-4o) for generation tasks—particularly in schema application. We chose o3-mini-high as our reasoning model due to its balance of accuracy and time efficiency. During early experiments, we found that it tended to generate overly profound outputs, so we reverted to GPT-4o for the generation steps. All prompts used in the system are available in Appendix A.
Next, we illustrate a typical interaction with the system through a walkthrough. In this scenario, the user wants to learn how to write an HCI paper abstract using a sample set of 15 CHI’24 best paper abstracts. The user also gathers the corresponding paper titles as input context.
3.1. Preprocessing Step for Multimodal Input
The user enters the system, which displays an example input node for specifying their goal and example set. To begin, the user enters the creation goal: “Write an HCI paper abstract”, and uploads a collection of 15 CHI’24 Best Paper abstracts along with the corresponding paper titles. The examples and corresponding inputs are stored in pairs within the system.
Although not applicable in this specific use case, the system is capable of analyzing multimodal examples. If multi-model content is detected, the system performs additional processing steps. For images, they are sent to GPT-4V to generate visual descriptions. For videos, the system extracts one frame per second as “screenshot” images and sends them to GPT-4V to generate visual descriptions and extract any available captions. The audio tracks of the videos are transcribed using Whisper. All parsed information is then combined into a structured textual file containing visual descriptions, captions (if available), and audio transcriptions.
Once the examples are processed and ready for analysis, the system generates a clustering operation node and prompts the user to click the “Run Clustering” button.
3.2. Stage 1: Clustering
Clustering (see Figure 3) identifies subclasses of examples and is crucial in preventing schema overgeneralization. When given a set of examples, it is often unclear if they share the same schema. Generalizing across structurally different examples can lead to weak and bland schemas. Therefore, we must perform clustering first to identify examples that share latent similarities with each other and then find the schema specific to each cluster.
After the user clicks the “Run Clustering” button, the system instructs the reasoning model to cluster the examples and provide reasoning. The system processes all of the examples simultaneously, producing and labeling clusters. The resulting clusters are then displayed as individual nodes on the Schemex interface.
For instance, Schemex maps the abstracts into 3 clusters: Empirical Studies (6 examples), Theoretical Contributions (4 examples) and System Design and Evaluation (5 examples).
The user can easily verify the clustering results by selecting a cluster node to view a side panel that explains the common features shared among its examples, including a matrix describing how closely each example aligns with these features. For example, Schemex identifies the common features of Theoretical Contributions papers to be: 1) Discussion of frameworks or concepts that advance understanding within the field, 2) Emphasis on implications for theory and policy, 3) Engagement with existing literature to draw new insights or reframe existing theories.
After checking the side panel, the user is satisfied with the clustering and decides to explore the Theoretical Contributions cluster in more depth. They click the ”Infer Dimensions” button on the cluster node.
3.3. Stage 2: Abstraction
Abstraction (see Figure 4) identifies structural commonalities among examples within each cluster. Unlike “features” that are used for clustering, the schemas formed during “abstraction” synthesize multiple examples into a general pattern of specific dimensions with descriptive attributes. It synthesizes multiple examples and extracts generalizable insights, a task that can often be overwhelming for individuals due to the high cognitive load involved.
Once the user clicks “Infer Dimensions” for the selected cluster, Schemex prompts the reasoning model to analyze the examples and extract the core elements or dimensions of the examples. For the Theoretical Contributions cluster, the key dimensions include: Contextual Framework , Gaps in Existing Literature, Methodological Approach, Findings and Contributions. The user can verify the generated schema by clicking on the dimension node to view further details about each dimension. This panel also includes an interactive Dimension-Example matrix that provides a visualization of how well each example conforms to each dimension. The matrix provides specific citations and justifications regarding whether each example fully, partially, or did not incorporate each dimension, enabling users to assess the system’s coherence by mapping its analysis back to the examples.
The next step is to “Infer Attributes”, including the dimension attributes, the key features within each dimension, and the overall attributes, the key descriptors of the full output. For example, for the Contextual Framework dimension, key attributes include a clearly defined population from which theoretical insights emerge or to which they apply, and the broader sociotechnical, cultural, or historical context that shapes the significance and applicability of the theory. For overall attributes, examples include a target word count range of 140–150 words and a narrative structure that presents the dimensions sequentially.
This approach distills executable guidelines for the specific cluster, ensuring that the output not only captures the key components (Dimensions) but also details what makes each component effective (Dimension-specific Attributes) and how to integrate these components into a cohesive output (Overall Attributes). At this point, the user has the first version of a usable schema.
3.4. Stage 3: Refinement via Contrasting Examples
The schema derived from Stage 2 is not perfect. For instance, while it specifies that important findings should be summarized, it does not clearly define how they should be written. To make the schema more precise and actionable, we employ an apply-and-test approach, iterating on the schema using contrastive refinement (see Figure 5).
Specifically, once the user has created an initial schema, they can click “Apply Schema” to prompt GPT-4o to generate Dimension values based on the current schema and an example’s input context (here is the paper title). This intermediate step allows users to 1) assess the accuracy of the identified dimensions and attributes before generating a final output 2) provide insight into the individual components behind the final product, deepening the user’s understanding of the system’s reasoning, as well as the task. Next, the user can generate a complete output (e.g., an HCI paper abstract) based on the generated dimension values and overall attributes. Note that the user can pick from a validation set (examples that came from the cluster was set separately and not used in the training) for schema application to check the schema effectiveness.
Schemex provides a color-coded interface that allows to user to easily compare the generated outputs with the original examples by dimension, highlighting potential improvements extracted by the reasoning model. These improvements are then used to iterate the schema. For example, the system might suggest that findings should be explicitly linked to existing literature and frameworks, rather than presented in isolation. After one iteration, the system generated a more detailed schema incorporating this improvement for the Findings dimension, such as adding the attribute “theoretical integration.” The iterated schema can then be used to generate a new set of dimension values and complete outputs. The cycle repeats until users are satisfied with the final schema. The user can keep exploring another cluster they are interested in.
Throughout their interaction with Schemex, users are involved in the process by evaluating whether the generated content meets their standards. By highlighting and citing content relevant to the AI’s reasoning decisions, Schemex encourages critical analysis of system output and constant comparison to the source material. These features support the user in their understanding of the AI and the paper abstract writing ask, and the original inputs, so they can more effectively act as a reflective evaluator.
Ultimately, users can utilize the resulting schemas as a guide to write their own HCI paper abstracts.
@@ 4.1. Data Preparation
To generate schema with o1-pro, we use the following prompt, which is equivalent to what we give Schemex as the task description and input. The prompt can be found in Appendix B. We ran models or pipelines in three conditions all once for each of the task. After getting all the schemas, we used GPT-4o to generate the outputs (as described in the System section “schema application” step). For each schema (of each cluster), we generated outputs for two randomly picked inputs from the cluster. Note that if a cluster just has one example, we just generated output for one input. Overall, we generated 20 schemas and 38 outputs for Condition 1, and 17 schemas and 33 outputs for Condition 2 and Condition 3, respectively.
@@ 4.2. Participants and Procedure
We invited twelve experts (two experts for each topic) to evaluate the schema and the accordingly outputs. The experts (average age 27.0, 7 female, 5 male) all have experience in the relevant domain and the specific task. For example, for news TikTok tasks, we invited journalism graduate students who had two years of experience in making news TikToks. Before the evaluation, we presented experts with examples of the schemas. We provided an evaluation rubric of the seven dimensions and guided them through it using practical examples. We requested them to read the original examples thoroughly before commencing ratings on schemas and outputs. For each of the tasks, two experts independently evaluated each dimension on a 7-point scale with justifications. When presenting the data to the experts, the condition information is removed, and the order of the condition shown to experts are counterbalanced between all the participants. The experts were compensated \$50/hour.
@@ 4.3. Evaluation Results and Findings
The interrater Agreement (within ±1 point) is 0.78, showing substantial agreement between the two experts’ judgment. This is deemed acceptable due to its highly subjective nature. We average the two experts’ scores for each schema and output. We run paired-sample Wilcoxon tests to compare the scores. See results in Table 1.
@@
Schemex Initial outperforms o1-pro in all seven dimensions, and it shows statistically significant improvements in three of the seven evaluation dimensions: D1: cluster quality (p=.046), D2: dimension and attribute quality (p=.026) and D5: usefulness of the schema (p=.032).
@@
For example, for the pitch deck test suite, one limitation of the Schemex full-version schema is that it places greater emphasis on specific storytelling patterns—such as step-by-step, roadmap-style narratives—which were prominent in the example set. While this leads to more precise alignment with those patterns, it may not generalize as well to pitches that follow less conventional structures. For instance, a founder who presents their mission, product, and traction in a fluid, vision-led narrative—without using explicit sequencing—might be harder to fully capture within the refined schema. The core logic is still there, but the schema may not recognize it as strongly. In contrast, the o1-pro schema is more focused on identifying the underlying reasoning—problem, solution, and evidence—regardless of how the story is told. This gives it more flexibility in handling diverse and novel formats, even when they diverge from patterns seen in the example set.
5. User Study
@@ 5.1. Protocol
We recruited eight participants (average age=26.6, four female, four male) for this study through a university mailing list and word-of-mouth. All participants reported being familiar or very familiar with ChatGPT. Each participant selected an example set of their choice. The example sets spanned diverse areas and modalities with 8-15 examples per set (see Table 2).
The participant was then given two tasks, (Task A) inducing a satisfactory schema using Schemex, and (Task B) inducing a satisfactory schema using o1-pro for the same example set. Half began with Task A, and half with Task B, with a brief demonstration before the Schemex task. Each task had a time limit of 30 minutes. After each task, participants were interviewed regarding their experience with the corresponding system. In these interviews, users rated questions (see questions in Section 5.2) on a 1-7 scale and provided reasons for their ratings. Each study session lasted approximately 80 minutes, with participants compensated \$25/hour. We run paired-sample Wilcoxon tests to compare the ratings.
@@ 5.2.3. RQ3 - Novel Reflection
Furthermore, exploration options provided by Schemex provided avenues for refinement they had not considered. For example, when using ChatGPT, P2 (Instagram Posts) had difficulty determining how to improve its outputs: “You have to know what to ask it in order to get what you want. It’s a little bit more frustrating.” He scored ChatGPT 2/7. P1 explains how the options provided by Schemex (scored 7/7) remedy this problem, citing the suggestions provided in the comparison step: “Schemex helps you with how you should iterate and provide you with an understanding of all of the opportunities.” He compares this to the experience using ChatGPT (scored 2/7): “I know what needs to change, but I don’t know what to change to. [I have to] act in a very corrective way.”
@@ 6.1. Supporting Schema Discovery through Interactive Structural Abstraction
The effectiveness of interactive schema discovery also depends on the capabilities of the underlying AI. Large language models—especially those optimized for reasoning—have recently become capable of meaningful structural analysis: clustering examples, identifying latent dimensions, and proposing abstractions that align with human interpretations. Schemex leverages these capabilities through a guided, multi-stage process: the AI suggests patterns, but always in a way that users can inspect and refine. This makes the system a collaborative partner rather than a black box, aligning well with best practices in human-AI interaction.
While Schemex was designed and evaluated for schema induction in creative and communicative tasks, the underlying approach is applicable to a wide range of domains. Many real-world problems involve extracting structure from complex, nonlinear data—for example, mapping workflows in organizational process mining, modeling user flows on websites, or tracing state transitions in software systems. These domains all require iterative exploration and abstraction to surface useful insights. The interactive structural abstraction model supported by Schemex provides a general framework for approaching such problems, suggesting promising directions for future applications.
@@ 6.3. Limitations and Future Work
Schemex currently supports multi-modal example analysis by converting images, videos, or audio into structured textual descriptions to enable LLM-based reasoning. While this approach allows schema induction across diverse content types, it limits the evaluation and application of schemas to a primarily textual modality. Future work could explore fully multi-modal interaction—for example, aligning schema elements with visual or interactive outputs—to help users better assess how schemas function across different media.
@@ 7. Conclusion
We introduced Schemex, an interactive system that supports schema induction through AI-assisted structural abstraction and contrastive refinement. By grounding abstract patterns in real examples and enabling iterative comparison, Schemex helps users uncover, evaluate, and refine the underlying structure of creative artifacts. Our study shows that Schemex improves user insight, confidence, and schema quality compared to a reasoning AI model baseline. This work highlights the potential of interactive, example-driven workflows to scaffold complex reasoning tasks, and opens new directions for supporting structure discovery across domains.

### v2

@@ 3.2. Motivating User Walkthrough
Sam clicks Infer Dimensions to dive deeper into the Theoretical Contributions cluster (see Figure 4). Schemex proposes four candidate dimensions: Contextual Framework, Gaps in Existing Literature, Methodological Approach, and Findings and Contributions. Each dimension is listed in the interface, and Sam opens the side panel to view how individual examples reflect it. A dimension-example matrix provides judgments for each dimension (yes/partial/no), again linked to citations from the abstracts.
Sam then clicks Infer Attributes. The system generates both dimension-specific attributes (e.g., Contextual Framework → Specific Population, Broader Context; Findings → Summary of Findings, Implications Discussed) and overall attributes (e.g., Word Count Range, Structured Format, Sequential Organization). The side panel shows another matrix, this time mapping attributes to examples, with highlighted spans that justify each label.
Sam notices that the “Specific Population” attribute feels too vague for her needs. She hovers over the label, edits it to “Focal Community,” and sees the change instantly reflected in the interface. She leaves other attributes unchanged, appreciating that the system allows lightweight edits. This process keeps abstractions tightly connected to evidence (DG1) while preserving Sam’s ability to refine them interactively (DG4).
3.2.3. Stage 3: Contrastive Refinement.
With an initial schema in place, Sam clicks Apply Schema to test its effectiveness (see Figure 5). Schemex first generates dimension values for each paper based on its title and dimension attributes. It then composes full draft abstracts that integrate both the dimension values and the overall attributes. The interface presents these drafts side by side with their corresponding gold examples. At this point, the system also applies the schema to the validation subset set aside earlier. Sentences are color-coded by dimension, allowing Sam to align the generated text with its counterpart in the original abstract.
During this process, Schemex detects gaps and inconsistencies between generated and gold examples. For instance, it flags that the most generated outputs summarize findings without linking them to existing theoretical frameworks. Based on this pattern, the system suggests adding a new attribute, theoretical integration, under the Findings dimension. Sam can then review such proposals with color-coded pairs, choosing to accept, reject, or edit them. Once reviewed, the schema updates to incorporate the approved suggestions, and the system regenerates outputs using the refined schema.
With each cycle of applying, inspecting, and refining, the schema becomes sharper and more transferable. This process makes schema quality visible through application (DG3) while keeping Sam in control (DG4).
Through this walkthrough, Sam gradually develops a schema that captures the structural essence of CHI theory paper abstracts. Each stage ensures that schemas are grounded in evidence (DG1), derived from structurally coherent clusters (DG2), tested and refined through practice (DG3), and shaped by active user oversight (DG4).
3.3. Key Design Features
3.3.1. Multimodal preprocessing.
To support schema induction over image and video examples, Schemex first converts multimodal artifacts into a unified textual representation before clustering and abstraction. A vision model generates visual descriptions of images and videos, while automatic speech recognition transcribes audio.
3.3.2. Stage 1: Clustering with rationales and matrices.
Schemex asks a reasoning model to analyze user-uploaded examples and propose clusters with rationales. The cluster side panel lists common features and presents a yes/partial/no example×feature matrix; clicking a cell brings up inline citations for that judgment. Users can scan quickly, change clusterings, and choose which cluster(s) to pursue next.
3.3.3. Stage 2: Abstraction into dimensions and attributes with evidence links.
Upon request, Schemex proposes dimensions, dimension-specific attributes, and overall attributes—each with definitions—and populates a dimension×example matrix with yes/partial/no judgments and citations. Users can edit, add, remove, or rename schema elements, or return to clustering if the emerging schema is too coarse or narrow. By organizing learning around relational structures and making alignment explicit, it helps novices articulate schemas while minimizing search or validation effort.
3.3.4. Stage 3: Contrastive refinement through apply-and-test.
Schemex operationalizes “test-and-revise” by applying the current schema. It produces per-dimension values, then composes a full draft under the overall attributes. The comparison view uses per-dimension colors to align generated text with similar sentences in the corpus and lists dimension-scoped improvement prompts. Users accept/reject prompts, modify the schema, and reapply to see effects.
3.3.5. Interaction scaffolds for novices.
A node-based canvas shows the pipeline and its artifacts; users can enter at any node to re-run or revise without losing context. A persistent side panel consolidates definitions, matrices, and inline citations so that evidence is always one click away. Localized controls—opening rationales, filtering matrices, accepting/rejecting suggestions, and editing schema elements—keep iteration fast while preserving agency.
3.4. Implementation
The frontend is implemented in React, using React Flow to render the node-based workflow and side-panel interactions. The backend is in Python, exposing FastAPI endpoints to orchestrate model calls and persist intermediate artifacts (with a Flask app). In this build, we use o3-mini-high for reasoning-intensive steps (cluster rationales; dimension/attribute inference) and GPT-4o for generation during schema application. For preprocessing, GPT-4V summarizes images/video frames and Whisper transcribes audio. Prompts used can be found in Appendix A.
@@ 4.1. Data Preparation
To generate schema with o1-pro, we use the prompt that is equivalent to what we give Schemex as the task description and input. The prompt can be found in Appendix B. We ran models or pipelines in three conditions all once for each of the task. After getting all the schemas, we used GPT-4o to generate the outputs (as described in the System section “schema application” step). For each schema (of each cluster), we generated outputs for two randomly picked inputs from the cluster. Note that if a cluster just has one example, we just generated output for one input. Overall, we generated 20 schemas and 38 outputs for Condition 1, and 17 schemas and 33 outputs for Condition 2 and Condition 3, respectively.
@@ 4.2. Participants and Procedure
We invited twelve experts (two experts for each topic) to evaluate the schema and the accordingly outputs. The experts (average age 27.0, 7 female, 5 male) all have experience in the relevant domain and the specific task. For example, for news TikTok tasks, we invited journalism graduate students who had at least two years of experience in making news TikToks; for the LinkedIn profile summary task, we invited university career coaches who regularly advise students on crafting professional profiles. Before the evaluation, we presented experts with examples of the schemas. We provided an evaluation rubric of the seven dimensions and guided them through it using practical examples. We requested them to read the original examples thoroughly before commencing ratings on schemas and outputs. For each of the tasks, two experts independently evaluated each dimension on a 7-point scale with justifications. When presenting the data to the experts, the condition information is removed, and the order of the condition shown to experts are counterbalanced between all the participants. The experts were compensated \$50/hour.
@@ 4.3. Evaluation Results and Findings
The interrater Agreement (within ±1 point) is 0.78, showing substantial agreement between the two experts’ judgments. This is deemed acceptable due to its highly subjective nature. We average the two experts’ scores for each schema and output. We run paired-sample Wilcoxon tests to compare the scores. See results in Table 1.
@@
Schemex Initial gets a higher average score than o1-pro in all seven dimensions, and it shows statistically significant improvements in three of the seven evaluation dimensions: D1: cluster quality (p=.046), D2: dimension and attribute quality (p=.026) and D5: usefulness of the schema (p=.032).
@@
For example, for the pitch deck test suite, one limitation of the Schemex full-version schema is that it places greater emphasis on specific storytelling patterns—such as step-by-step, roadmap-style narratives—which were prominent in the example set. While this leads to more precise alignment with those patterns, it may not generalize as well to pitches that follow less conventional structures. For instance, a founder who presents their mission, product, and market traction in a fluid, vision-led narrative—without using explicit sequencing—might be harder to fully capture within the refined schema. The core logic is still there, but the schema may not recognize it as strongly. In contrast, the o1-pro schema is more focused on identifying the underlying reasoning—problem, solution, and evidence—regardless of how the story is told. This gives it more flexibility in handling diverse and novel formats, even when they diverge from patterns seen in the example set. These limitations could be mitigated in practice through human oversight, as well as by incorporating a more diverse range of examples or adjusting the clustering granularity to a coarser level.
5. User Study
@@ 5.1. Protocol
We recruited eight participants (average age=26.6, four female, four male) through a university mailing list and word-of-mouth. All participants reported being familiar or very familiar with ChatGPT. Each participant selected an example set of their choice. The example sets spanned diverse areas and modalities with 8-15 examples per set (see Table 2).
The participant was then given two tasks, (Task A) inducing a satisfactory schema using Schemex, and (Task B) inducing a satisfactory schema using o1-pro for the same example set. Half began with Task A, and half with Task B, with a brief demonstration before the Schemex task. Each task had a time limit of 30 minutes. After each task, participants were interviewed regarding their experience with the corresponding system. In these interviews, users rated questions (see questions in Section 5.2) on a 1-7 scale and provided reasons for their ratings.
Each study session lasted approximately 80 minutes, with participants compensated \$25/hour. We run paired-sample Wilcoxon tests to compare the ratings.
@@ 5.2.3. RQ3 - Novel Reflection
Furthermore, exploration options provided by Schemex provided avenues for refinement they had not considered. For example, when using ChatGPT, P2 (Instagram Posts) had difficulty determining how to improve its outputs: “You have to know what to ask it in order to get what you want. It’s a little bit more frustrating.” He scored ChatGPT 2/7. P1 explains how the options provided by Schemex (scored 7/7) remedy this problem, citing the suggestions provided in the comparison step: “Schemex helps you with how you should iterate and provide you with an understanding of all of the opportunities.” She compares this to the experience using ChatGPT (scored 2/7): “I know what needs to change, but I don’t know what to change to. [I have to] act in a very corrective way.”
@@ 6.1. Supporting Schema Discovery through Interactive Structural Abstraction
The effectiveness of interactive schema discovery also depends on the capabilities of the underlying AI. Large language models—especially those optimized for reasoning—have recently become capable of meaningful structural analysis: clustering examples, identifying latent dimensions, and proposing abstractions that align with human interpretations. Schemex leverages these capabilities through a guided, multi-stage process: the AI suggests patterns, but always in a way that users can inspect and refine. This makes the system a collaborative partner rather than a black box, aligning well with recommended practices in human-AI interaction.
While Schemex was designed and evaluated for schema induction in creative and communicative tasks, the underlying approach is applicable to a wide range of domains. Many real-world problems involve extracting structure from complex, nonlinear data—for example, mapping workflows in organizational process mining, modeling user flows on websites, or tracing state transitions in software systems. These domains all require iterative exploration and abstraction to surface useful insights. The interactive structural abstraction model supported by Schemex provides a framework for approaching such problems, suggesting promising directions for future applications.
@@ 6.3. Limitations and Future Work
Schemex currently supports multi-modal example analysis by converting images, videos, or audio into structured textual descriptions to enable LLM-based reasoning. While this approach allows schema induction across diverse content types, it limits the evaluation and application of schemas to a primarily textual modality. Future work could explore fully multi-modal comparison and interaction—for example, aligning schema elements with visual or interactive outputs—to help users better assess how schemas function across different media.
@@ 7. Conclusion
We introduced Schemex, an interactive system that supports schema induction through AI-assisted structural abstraction and contrastive refinement. By grounding abstract patterns in real examples and enabling iterative comparison, Schemex helps users uncover, evaluate, and refine the underlying structure of creative artifacts. Our study shows that Schemex significantly improves user insight, reflection, and schema quality compared to a reasoning AI model baseline. This work highlights the potential of interactive, example-driven workflows to scaffold complex reasoning tasks, and opens new directions for supporting structure discovery across domains.